Micah is 101 years old. Alma is 44 years old. How many years ago was Micah’s age 4 times Alma’s age?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

Much like another one recently asked and answered.

See this one for how:
https://www.algebra.com/algebra/homework/word/age/Age_Word_Problems.faq.question.1125349.html
https://www.algebra.com/algebra/homework/word/age/Age_Word_Problems.faq.question.1125349.html

Answer by ikleyn(52788)

.
Let x= "how many years ago . . . ".


Then the equation is


    101 - x = 4(44-x)


    101 - x = 176 - 4x


    4x - x = 176 - 101

    3x     =  75    ====>  x = 75/3 = 25.


Answer.  25 years ago.

You can check on your own that the solution is correct.

------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.